A. Molteni Motors Inc. recently reported $2.25 million of net income. Its EBIT was $5.25 million, and its tax rate was 40%. What was its interest expense? (Hint: Write out the headings for an income statement and then fill in the known values. Then divide $2.25 million net income by 1 − T = 0.6 to find the pre-tax income. The difference between EBIT and taxable income must be the interest expense.) Round your answer to the nearest dollar. Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000.

B.

Corporate Tax Liability

To complete the assignments listed below, refer to the Table 2-1.

The Talley Corporation had a taxable income of $350,000 from operations after all operating costs but before (1) interest charges of $70,000, (2) dividends received of $14,000, (3) dividends paid of $17,500, and (4) income taxes.

What are the firm's income tax liability and its after-tax income? Round your answers to two decimal places.

Income tax liability $  
After-tax income $  

What are the company's marginal and average tax rates on taxable income? Round your answers to two decimal places.

Marginal tax rate %
Average tax rate %
Answer #1

A)

We have the following info:

EBIT 5.25
Interest x
Tax 0.40*(5.25-x)
NI 2.25

Tax = PBT*0.40

PBT=5.25-x

Hence 5.25-x-(2.1-0.4x)=2.25

5.25-x-2.1+0.4x=2.25

3-x-2.1=-0.4x

0.9-x=-0.4x

0.9=0.6x

x=1.5

hence interest =$1.5 million or $1,500,000

For Part B information missing as to tax rates